# SaVe-Tag: Semantic-aware Vicinal Risk Minimization for Long-Tailed Text-Attributed Graphs

### v3

@@ Abstract
Real-world graph data often follows long-tailed distributions, making it difficult for Graph Neural Networks (GNNs) to generalize well across both head and tail classes. Recent advances in Vicinal Risk Minimization (VRM) have shown promise in mitigating class imbalance with numeric interpolation; however, existing approaches largely rely on embedding-space arithmetic, which fails to capture the rich semantics inherent in text-attributed graphs. In this work, we propose our method SaVe-Tag (Semantic-aware Vicinal Risk Minimization for Long-Tailed Text-Attributed Graphs), a novel VRM framework that leverages Large Language Models (LLMs) to perform text-level interpolation, generating on-manifold, boundary-enriching synthetic samples for minority classes. To mitigate the risk of noisy generation, we introduce a confidence-based edge assignment mechanism that uses graph topology as a natural filter to ensure structural consistency. We provide theoretical justification for our method and conduct extensive experiments on benchmark datasets, showing that our approach consistently outperforms both numeric interpolation and prior long-tailed node classification baselines. Our results highlight the importance of integrating semantic and structural signals for balanced and effective learning on text-attributed graphs. The source code is available here: https://github.com/LWang-Laura/SaVe-TAG.
Figure 1: Comparison of numeric and LLM-based interpolation.(a) Illustration of the two interpolation strategies; (b) Advantages of LLM-based interpolation in producing boundary-enriching samples are highlighted (visualization uses synthetic data for illustration); (c) UMAP projection of the original and interpolated samples from the Cora dataset—the original data boundary is outlined, with numeric samples mostly staying within and LLM-generated samples extending beyond.
1 Introduction
Graph Neural Networks (GNNs) have proven effective for node classification by modeling structural dependencies in graph data. However, real-world graphs frequently exhibit long-tail distributions [1, 2, 3], making it difficult to learn from underrepresented classes.
@@ 1 Introduction
Recent approaches to data imbalance often adopt the Vicinal Risk Minimization (VRM) principle [4], which minimizes loss over a synthetic neighborhood of the training data—extended from Empirical Risk Minimization (ERM) [5] that focuses solely on the existing data points. Typically, VRM is instantiated via interpolation techniques like Mixup [6] and SMOTE [7], and has been adapted for long-tailed graph learning in methods including GraphSMOTE [8] and MixupForGraph [9], etc. However, these methods often apply mathematic operations over input embeddings or hidden features, failing to preserve the rich manifold structure of text-attributed graphs (TAGs).
Figure 2: Effect of VRM.
@@
Advances in large language models (LLMs) present new opportunities for data augmentation by offering external knowledge and creative generation. However, prior LLM-based augmentation approaches typically rely on zero-shot or few-shot prompting [10, 11], which usually fail to generate samples within the neighborhood of minority instances, as required by the VRM principle. Nevertheless, VRM has been demonstrated critical in improving long-tailed node classification. As shown in Figure 2a, VRM-guided numeric interpolation with nearest neighbors yields synthetic embeddings that outperform ERM-style embedding duplication.
@@
Motivated by this insight, we hypothesize that prompting strategies explicitly designed to mimic interpolation will surpass standard zero/few-shot approaches, which is subsequently validated by the results in Figure 2b. Furthermore, integrating controlled prompting into LLM generation can alleviate common pitfalls of numeric interpolation, enabling manifold-aware, boundary-enriching augmentation that effectively exploits the knowledge encoded in LLMs (see Figure 1).
@@
In summary, we introduce a novel interpolation framework, SaVe-Tag, which unifies textual semantics and graph topology. Our method employs LLM-based interpolation to expand the vicinal boundary and reduce vicinal risk, followed by a topology-aware filtering mechanism that preserves alignment within the underlying data distribution. Our key contributions are summarized as follows:
  • To the best of our knowledge, we present the first semantic-aware VRM framework that employs LLM-based interpolation for node classification on long-tailed text-attributed graphs (TAGs).
@@
  • We conduct extensive experiments demonstrating that our method outperforms embedding-based and long-tail graph baselines, highlighting the value of topology-aware filters and semantic benefits from LLMs.
2 Problem Formulation
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{T})$ be an attributed graph with $|\mathcal{V}|=n$ nodes, feature matrix [^1] For text-attributed graphs, each row of $\mathcal{X}$ stores a tokenized document. $\mathcal{X}\in\mathbb{R}^{n\times d}$, and labels $\mathcal{Y}=\{y_{v}|v\in\mathcal{V}\}$ with $y_{v}\in\!\{1,\dots,C\}$. Class imbalance induces a skewed empirical distribution $\mathcal{P}$ on $(\mathcal{X},\mathcal{Y})$. Our goal is to learn a node-level classifier $f_{\theta}:\mathbb{R}^{d}\!\!\times\!\mathcal{V}\!\to\!\Delta^{C-1}$, e.g., a GNN, that maximizes classification performance.
Figure 3: An overview of SaVe-Tag. Given an input graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{T})$, we perform LLM-based interpolation to synthesize boundary-enriching samples $\hat{t}$ that minimize vicinal risk. A confidence function is then pre-trained on the original graph and later used to assign edges to the synthetic nodes via a top-$k$ selection strategy. When the resulting augmented graph $\mathcal{G}^{\prime}=(\mathcal{V}^{\prime},\mathcal{E}^{\prime},\mathcal{T}^{\prime})$ is processed by a GNN, such edge assignment incorporates well-aligned nodes into their vicinity while isolating noisy samples.
3 SaVe-Tag Framework
An overview of our framework is shown in Figure 3, with details in Algorithm 1 from Appendix B.
@@ 3 SaVe-Tag Framework
Let $\mathcal{C}_{t}\subseteq\{1,\dots,C\}$ denote the set of tail (minority) classes. For each node $v_{1}\in\mathcal{V}$ with label $y_{v_{1}}\in\mathcal{C}_{t}$, we define its set of vicinal twins as
$\mathcal{N}_{k}^{\text{ (tail)}}(v_{1}):=\bigl{\{}v_{2}|\,v_{2}\in\mathcal{V}\;,\;y_{v_{2}}=y_{v_{1}},\;v_{2}\in\mathcal{N}_{k}(v_{1})\bigr{\}},$
i.e., the nodes from the same tail class that lie within the $k$-hop (or $k$-nearest) neighborhood $\mathcal{N}_{k}(v_{1})$ of $v_{1}$. As $(v_{1},v_{2})$ are in the vicinity, synthesizing new nodes from them follows the VRM principle.
LLM-based interpolation.
For any arbitrary pair of vicinal twins $(v_{i},v_{j})\in\mathcal{P}$, we query a temperature-controlled, class-conditioned LLM $\mathcal{L}_{\tau}$ to create $\hat{t}_{ij}\;=\;\mathcal{L}_{\tau}\!\bigl{(}t_{v_{i}},\,t_{v_{j}}\,\bigm{|}\,y_{v_{i}}\bigr{)}$ for $\hat{v}_{ij}$ such that $y_{\hat{v}_{ij}}=y_{v_{i}}$ and collect the new nodes in $\bar{\mathcal{V}}:=\{\hat{v}_{ij}\}$ with texts $\bar{\mathcal{T}}:=\{\hat{t}_{ij}\}$.
Edge Assignment with Confidence Function.
A frozen encoder $\phi:\mathcal{T}\!\to\!\mathbb{R}^{d^{\prime}}$ produces $z_{v}=\phi(t_{v}),\;z_{\hat{v}}=\phi(\hat{t})$, while a confidence function $\kappa:[0,1]$ is trained on the original graph. Each $\hat{v}$ connects to the $k$ highest-scoring neighbors
$\mathcal{N}_{k}(\hat{v})=\operatorname*{arg\,top}_{k,\;u\in\mathcal{V}}\bigl{[}\kappa_{\omega}(z_{u})\,\langle z_{\hat{v}},z_{u}\rangle\bigr{]},\qquad\bar{\mathcal{E}}_{\hat{v}}=\{(\hat{v},u)\mid u\in\mathcal{N}_{k}(\hat{v})\},$
and is identified as isolated if $\max_{u}\kappa_{\omega}(z_{u})<\tau$. Setting $\mathcal{V}^{\prime}=\mathcal{V}\cup\bar{\mathcal{V}},\;\mathcal{E}^{\prime}=\mathcal{E}\cup\!\!~{}\bigcup_{\hat{v}}\bar{\mathcal{E}}_{\hat{v}},\;\mathcal{T}^{\prime}=\mathcal{T}\cup\bar{\mathcal{T}}$ yields $\mathcal{G}^{\prime}=(\mathcal{V}^{\prime},\mathcal{E}^{\prime},\mathcal{T}^{\prime})$, on which the GNN $f_{\theta}$ is further trained.
4 SaVe-Tag Theory
4.1 LLM-based Interpolation
In this subsection, we formalize how class-conditioned generation preserves semantic consistency, enriches boundary regions, and improves generalization through Vicinal Risk Minimization.
4.1.1 Setup and Notation
Embeddings and class manifolds.
Let $\phi:\mathcal{X}\to\mathbb{R}^{d^{\prime}}\!,~{}z_{v}\;:=\;\phi(x_{v})$ denote the penultimate embedding produced by $f_{\theta}$. For each class $c$ we write $\mathcal{M}_{c}=\bigl{\{}z_{v}:y_{v}=c\bigr{\}}$ for the (empirical) class manifold in feature space and set $\mathcal{M}:=\bigcup_{c=1}^{C}\mathcal{M}_{c}$.
Text generator.
We use $\mathcal{L}_{\tau}$ to denote the LLM text generator with temperature $\tau$. Given seeds $t_{1},t_{2}\!\in\!\mathcal{T}$ (raw texts) and an explicit class prompt $c$, the generator returns $\hat{t}\;=\;\mathcal{L}_{\tau}(t_{1},t_{2}\mid c).\;$
4.1.2 Manifold-preserving Interpolation
@@
Let $\phi(t_{1}),\phi(t_{2})\in\mathcal{M}_{c}$. Under the linearity constraint, numeric interpolation yields the combination based on convex assumption [12]:
$z_{\lambda}\;=\;\lambda\,\phi(t_{1})\;+\;(1-\lambda)\,\phi(t_{2})~{}\text{ for}~{}\lambda\in(0,1)$
If the class manifold $\mathcal{M}_{c}$ is non-convex, then ${}~{}z_{\lambda}\notin\mathcal{M}_{c}~{}~{}\text{ for all }\lambda\in(0,1)~{}$ [13, 14].
@@
Let $t_{1},t_{2}\in\mathcal{T}$ satisfy ${}\phi(t_{1}),\;\phi(t_{2})\;\in\;\mathcal{M}_{c}~{}$, and $\hat{t}=\mathcal{L}_{\tau}\bigl{(}t_{1},\,t_{2}\mid c\bigr{)}$ condition on class prompt $c$
Then there exists a small $\delta>0$ such that, with probability at least  $1-\delta$ , $\phi(\hat{t})\;\in\;\mathcal{M}_{c}~{},$
i.e. the LLM preserves class consistency and manifold under label-guided generation:

### v5

# SaVe-TAG: LLM-based Interpolation for Long-Tailed Text-Attributed Graphs

@@ Abstract.
Real-world graph data often follows long-tailed distributions, making it difficult for Graph Neural Networks (GNNs) to generalize well across both head and tail classes. Recent advances in Vicinal Risk Minimization (VRM) have shown promise in mitigating class imbalance with numeric interpolation; however, existing approaches largely rely on embedding-space arithmetic, which fails to capture the rich semantics inherent in text-attributed graphs. In this work, we propose our method SaVe-TAG (Semantic-aware Vicinal Risk Minimization for Long-Tailed Text-Attributed Graphs), a novel VRM framework that leverages Large Language Models (LLMs) to perform text-level interpolation, generating on-manifold, boundary-enriching synthetic samples for minority classes. To mitigate the risk of noisy generation, we introduce a confidence-based edge assignment mechanism that uses graph topology as a natural filter to ensure structural consistency. We provide theoretical justification for our method and conduct extensive experiments on benchmark datasets, showing that our approach consistently outperforms both numeric interpolation and prior long-tailed node classification baselines. Our results highlight the importance of integrating semantic and structural signals for balanced and effective learning on text-attributed graphs. The source code is publicly available at: https://github.com/LWang-Laura/SaVe-TAG.
Figure 1. Motivation. (a) Numeric vs. LLM-based interpolation.  (b) LLM-based interpolation enriches decision boundaries (on synthetic data).  (c) UMAP of original and interpolated samples on Cora; numeric interpolations remain mostly within class boundaries, while LLM-based samples extend beyond.  (d) Vicinal Risk Minimization (VRM) outperforms Empirical Risk Minimization (ERM) in data augmentation.
LLM Data Augmentation, Text-Attributed Graphs, Class Imbalance
journalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Republic of Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770854.3780311isbn: 979-8-4007-2258-5/2026/08ccs: Computing methodologies Machine learning
1. Introduction
@@
Graph Neural Networks (GNNs) have proven effective for node classification by modeling structural dependencies (Kipf and Welling, 2016; Hamilton et al., 2017; Rong et al., 2020). However, real-world graphs frequently exhibit long-tail distributions (Ma et al., 2025; Wang et al., 2021b, 2024), making it difficult to learn underrepresented classes. Recent approaches to data imbalance often adopt the Vicinal Risk Minimization (VRM) principle (Chapelle et al., 2000b), which minimizes loss over a synthetic neighborhood of the training data—extending Empirical Risk Minimization (ERM) that focuses solely on the existing data (Vapnik and Vapnik, 1998). Typically, VRM is instantiated via interpolation like Mixup (Zhang et al., 2017) and SMOTE (Chawla et al., 2002), and has been adapted for long-tailed graph learning (Zhao et al., 2021; Wang et al., 2021a). However, these methods often apply mathematic operations over input embeddings or hidden features, failing to preserve the rich manifold structure of text-attributed graphs (TAGs).
@@
Advances in large language models (LLMs) present new opportunities for data augmentation by offering external knowledge and creative generation. However, existing work on LLM-based augmentation typically rely on zero-shot or few-shot prompting (Li et al., 2023; Yu et al., 2024), which usually fail to generate samples within the neighborhood of minority instances, as required by VRM principle. Nevertheless, VRM has been shown critical in data augmentation, especially in improving long-tailed node classification. As seen in Figure 1d (i), VRM-guided numeric interpolation with nearest neighbors outperform ERM-style embedding duplication in node classification.
@@
Motivated by this insight, we hypothesize that prompting strategies explicitly designed to mimic numeric interpolation will surpass standard zero/few-shot approaches, validated by the results in Figure 1d (ii). Furthermore, integrating controlled prompting into LLM generation can alleviate common pitfalls of numeric interpolation, enabling manifold-aware, boundary-enriching augmentation that effectively exploits the knowledge encoded in LLMs (see Figure 1).
@@
In summary, we introduce a novel interpolation framework, SaVe-TAG, which unifies textual semantics and graph topology in data augmentation. Our method employs LLM-based interpolation to expand the vicinal boundary and reduce vicinal risk, followed by a topology-aware filtering mechanism that preserves alignment within the underlying data distribution. Our key contributions are:
  • To the best of our knowledge, we present the first semantic-aware VRM framework that employs LLM-based interpolation for node classification on long-tailed text-attributed graphs.
@@
  • We conduct experiments showing our method outperforms embedding-based augmentations and long-tail graph baselines, driven by topology-aware filters and LLM semantic benefits.
00footnotetext: The source code: https://github.com/LWang-Laura/SaVe-TAG.
Figure 2. An overview of SaVe-TAG. Given an input graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{T})$, we perform LLM-based interpolation on identified vicinal twins to synthesize boundary-enriching samples $\hat{t}$ that minimize vicinal risk. A confidence function is then pre-trained on the original graph and later used to assign edges to the synthetic nodes via a top-$k$ selection strategy. When the resulting augmented graph $\mathcal{G}^{\prime}=(\mathcal{V}^{\prime},\mathcal{E}^{\prime},\mathcal{T}^{\prime})$ is processed by a GNN, such edge assignment incorporates well-aligned nodes into their vicinity while isolating noisy samples.
2. Problem Formulation
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{T})$ be an attributed graph with node set $\mathcal{V}$, edge set $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$, and per‑node texts $\mathcal{T}=\{t_{v}\}_{v\in\mathcal{V}}$. Let $\mathcal{X}$ be the feature matrix, where each row corresponds to a tokenized text $t_{v}\in\mathcal{T}$. Each node $v$ is associated with a label $y_{v}\in\{1,\dots,C\}$, where $C$ is the total number of classes, forming the label set $\mathcal{Y}=\{\;y_{v}\mid v\in\mathcal{V}\;\}$. Class imbalance induces a skewed empirical distribution on $(\mathcal{X},\mathcal{Y})$. Our goal is to learn a node-level classifier $f_{\theta}$ (e.g., a GNN), that maximizes classification performance.
3. SaVe-TAG
In this section, we present the methodology of SaVe-TAG. We begin with an overview of the main pipeline ($\S$ 3.1), followed by detailed descriptions of its two core components: LLM-based Vicinal Risk Minimization ($\S$ 3.2) and Graph as Natural Filters ($\S$ 3.3).
3.1. Overview
The overall pipeline of SaVe-TAG is illustrated in Figure 2, with the corresponding pseudocode provided in Algorithm 1 (Appendix B). Specifically, it comprises the following steps:
  1. (1)
    Vicinal Twin Identification: identifying same-label vicinal twins in tail classes. For each tail-class node $v$, find $k$-NN same-label neighbors, denoted as $\mathcal{N}_{k}^{(\mathrm{tail})}(v)$.
  2. (2)
    LLM Synthesis: Synthesize class-consistent texts from vicinal twin pairs using a class-conditioned LLM. Specifically, given node texts $(t_{v_{i}},t_{v_{j}})$ from a vicinal twin pair $(v_{i},v_{j})$, we query the LLM at temperature $\tau$ (denoted as $\mathcal{L}_{\tau}$) to generate a synthetic text $\hat{t}$. The resulting synthetic node $\hat{v}$ is assigned the same label as the original twin, i.e., $y_{\hat{v}}=y_{v_{i}}$.
  3. (3)
    Confidence-aware Edge Assignment: Assign confidence-aware edges that pull well-aligned samples toward the correct manifold while isolating noisy ones. All texts are encoded using a frozen encoder $\phi$. For each newly generated text-attributed node $\hat{v}$, we connect it to high-confidence neighbors identified via a learned scoring function $\kappa(\cdot)$ for confidence.
  4. (4)
    GNN Training: Train a message-passing GNN on the augmented graph $\mathcal{G}^{\prime}=(\mathcal{V}^{\prime},\mathcal{E}^{\prime},\mathcal{T}^{\prime})$, where $\mathcal{V}^{\prime}$ is the augmented node set, $\mathcal{E}^{\prime}$ is the augmented edge set, and $\mathcal{T}^{\prime}=\{\;t_{v}\mid v\in\mathcal{V}^{\prime}\;\}$ contains the corresponding node texts of $\mathcal{V}^{\prime}$.
Steps (1) and (2) are incorporated into LLM-based Vicinal Risk Minimization ($\S$ 3.2), while steps (3) and (4) are unified under Graph as Natural Filters ($\S$ 3.3), detailed below.
3.2. LLM‑based Vicinal Risk Minimization
Here, we formalize how class-conditioned generation preserves semantic manifolds, densifies boundary regions, and improves generalization through Vicinal Risk Minimization (VRM), and present theoretical analyses of these properties. Complete proofs of all theorems are provided in Appendix D.
@@ 3.2. LLM‑based Vicinal Risk Minimization
Let $\mathcal{C}_{t}\subseteq\{1,\ldots,C\}$ denote the set of tail classes. For any text-attributed node $v_{1}$ with $y_{v_{1}}\in\mathcal{C}_{t}$, we define its vicinal twins as
(1) $\mathcal{N}_{k}^{(\mathrm{tail})}(v_{1}):=\bigl\{\,v_{2}\in\mathcal{V}\;\bigm|\;y_{v_{2}}=y_{v_{1}},\,v_{2}\in\mathcal{N}_{k}(v_{1})\bigr\},$
that is, same-label neighbors within the $k$-NN vicinity of $v_{1}$. Synthesizing new nodes from such pairs aligns with the VRM principle, as $(v_{1},v_{2})$ are close in both label and representation. We refer to such pairs as vicinal twins.
LLM‑based interpolation.
For any pair of vicinal twins $(v_{i},v_{j})$ and their corresponding text attributes $t_{v_{i}}$ and $t_{v_{j}}$, we generate a synthetic text $\hat{t}_{ij}$ by querying the LLM ( denoted as $\mathcal{L}_{\tau}$): $\hat{t}_{ij}=\mathcal{L}_{\tau}(t_{v_{i}},t_{v_{j}}\mid y_{v_{i}})$. This yields a new text-attributed node $\hat{v}_{ij}$ with label $y_{\hat{v}_{ij}}=y_{v_{i}}$. We collect the set of synthetic nodes as $\bar{\mathcal{V}}:={\hat{v}_{ij}}$ and their corresponding texts as $\bar{\mathcal{T}}:={\hat{t}_{ij}}$.
3.2.1. Staying on the Manifold
Let $\phi:\mathcal{T}\to\mathcal{X}$ denote a frozen text encoder, and write $x_{v}:=\phi(t_{v})$ for the embedding of node $v$ with text $t_{v}\in\mathcal{T}$. For each class $c\in\{1,\dots,C\}$, define the class manifold in feature space by
$\mathcal{M}_{c}:=\{\,x_{v}\mid y_{v}=c\,\},\qquad\mathcal{M}:=\bigcup_{c=1}^{C}\mathcal{M}_{c}.$
Let $(v_{1},v_{2})$ be a same‑label pair (vicinal twins) with $y_{v_{1}}=y_{v_{2}}=c$ and corresponding texts $(t_{1},t_{2})$.
@@ 3.2.1. Staying on the Manifold
If $\mathcal{M}_{c}$ is non-convex, there exist $t_{1},t_{2}$ with $\phi(t_{1}),\phi(t_{2})\in\mathcal{M}_{c}$ and $\lambda\in(0,1)$ such that $x_{\lambda}:=\lambda\,\phi(t_{1})+(1{-}\lambda)\,\phi(t_{2})\notin\mathcal{M}_{c}$ (Zhang et al., 2018; Guo et al., 2019; Baena et al., 2022).
Interpretation. Purely numeric mixup can cross gaps between modes in a curved or multi‑modal class manifold—common in text semantics—yielding ambiguous training targets.
@@
Let a synthetic text $\hat{t}\sim\mathcal{L}_{\tau}(t_{1},t_{2}\mid c)$ be a sample from an LLM conditioned on class $c$.
Suppose the class‑conditional distribution $p\;(\;t\mid y{=}c\;)$ satisfies $\Pr_{t\sim p(\cdot\mid y=c)}\;[\phi(t)\in\mathcal{M}_{c}]\geq 1-\delta$ (where $\Pr$ denotes probability) for some small error $\delta\in[0,1)$, and that $\mathcal{L}_{\tau}(\cdot\mid c)$ approximates $p(\cdot\mid y{=}c)$ with total‑variation error at most $\epsilon$. Then the generated text $\hat{t}$ will be mapped to the class-$c$ manifold with probability at least $1-(\delta+\epsilon)$, denoted as $\Pr\big[\phi(\hat{t})\in\mathcal{M}_{c}\big]\;\geq\;1-(\delta+\epsilon).$
Interpretation. Conditioning the generator on class labels ensures that most synthetic texts remain on the corresponding class manifold, motivating class‑conditioned prompting in SaVe-TAG.
3.2.2. Boundary‑Enriching Effect of LLM Samples
Let $\mathcal{X}\subseteq\mathbb{R}^{d}$ be the input space and $\mathcal{Y}=\{1,\dots,C\}$ the label set. A classifier $f$ outputs logits $f^{(c)}(x)$ for class $c\in\mathcal{Y}$ and the logit margin is denoted as $\gamma$. We augment $\mathcal{X}$ with $m$ LLM‑generated pairs $\{(\hat{x}_{j},\hat{y}_{j})\}_{j=1}^{m}$ to form ${\mathcal{X}}^{\prime}:=\mathcal{X}\cup\{(\hat{x}_{j},\hat{y}_{j})\}_{j=1}^{m}$.
Definition 3.3 (The Minimum Margin).
For a labeled example $(x,y)$, define its (logit) margin $\gamma(x,y)$ to be the difference between the model’s score for the true class $y$ and the highest score assigned to any other class, denoted as: $\gamma(x,y)\;:=\;f^{(y)}(x)\;-\;\max_{c\neq y}f^{(c)}(x),$
and for any labeled set $S$, define the minimum margin $\gamma_{\min}(S)\;:=\min_{(x,y)\in S}\gamma(x,y)$.
Definition 3.4 (Boundary‑Coverage Rate (BCR)).
A labeled sample $(x,y)$ is a boundary sample if the majority of its $k$ Nearest Neighbors have labels different from $y$ (Han et al., 2005). Boundary‑Coverage Rate (BCR) is the fraction of boundary samples in $\mathcal{X}^{\prime}$ .
Theorem 3.5 (Margin Lower Bound).
Let $\eta>0$ be a fixed margin slack parameter,  $\gamma_{\min}(\mathcal{X})$ denote the minimum margin achieved by a model trained on the original dataset $\mathcal{X}$, and $\gamma_{\min}(\mathcal{X}^{\prime})$ denote the minimum margin after retraining on the augmented dataset $\mathcal{X}^{\prime}$.
Then, adapting the argument from (Oh and Yun, 2023), the minimum margin after augmentation and retraining decreases by at most $\eta$ times the fraction of non-boundary samples, that is:
$\gamma_{\min}\bigl(\mathcal{X}^{\prime}\bigr)\;\geq\;\gamma_{\min}(\mathcal{X})\;-\;\eta\,\bigl(1-\mathrm{BCR}\bigr).$
Interpretation. Larger $\mathrm{BCR}$ means more near‑boundary samples; the bound shows the margin drop is capped by $\delta(1-\mathrm{BCR})$, so higher $\mathrm{BCR}$ improves robustness near decision boundaries (Han et al., 2005).
3.2.3. Vicinal Risk Minimization with LLM-based Neighborhood
Following VRM (Chapelle et al., 2000b), we define, for each training text $t_{i}$ with class label $c_{i}\!\in\!\{1,\dots,C\}$, an LLM‑based vicinal distribution
$\mathcal{V}_{\mathcal{L}_{\tau}}(t_{i})\;=\;\sum_{j=1}^{n}w_{ij}\,\Pr\;[\;\hat{t}=\mathcal{L}_{\tau}(t_{i},t_{j}\mid c_{i})\;],\quad w_{ij}\!\geq\!0,\;\;\sum_{j}w_{ij}=1,$
The corresponding vicinal risk is
$R_{\mathrm{vrm}}(f_{\theta})\;=\;\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\hat{t}\sim\mathcal{V}_{\mathcal{L}_{\tau}}(t_{i})}\!\left[\,\mathcal{J}\;\!\big(f_{\theta}(\phi(\hat{t})),\,c_{i}\big)\right],$
where $\mathbb{E}$ is the expectation operator, $\phi$ is a (frozen) text encoder, $f_{\theta}$ is the classifier, and $\mathcal{J}$ is the per‑example loss (e.g., cross‑entropy).
Theorem 3.6 (On‑manifold vicinal risk).
If Theorem 3.2 holds, then $R_{\mathrm{vrm}}(f_{\theta})$ is evaluated on‑manifold with probability at least $1-\delta$ (over the randomness of LLM sampling).
Interpretation. Thm. 3.6 ensures VRM mostly trains on valid (on‑manifold) synthetic data.
Theorem 3.7 (Boundary coverage $\Rightarrow$ lower vicinal risk).
Let $f_{\theta}^{\mathrm{aug}}$ denote the model trained by VRM on the augmented dataset $\mathcal{X}^{\prime}$, and let $f_{\theta}^{\mathrm{orig}}$ denote the model trained without augmentation on $\mathcal{X}$. Assume the loss function $\mathcal{J}$ is $L$-Lipschitz in its first argument. Then, for any tolerance parameter $\zeta>0$,
$R_{\mathrm{vrm}}(f_{\theta}^{\mathrm{aug}})\;\leq\;R_{\mathrm{vrm}}(f_{\theta}^{\mathrm{orig}})\;-\;L\,\gamma_{0}\,(\mathrm{BCR}-\zeta)\;+\;\mathcal{O}(\delta),$
where $L$ is the Lipschitz constant of the loss (i.e., its sensitivity to input changes), $\gamma_{0}:=\gamma_{\min}(\mathcal{X})$ is the minimum margin before augmentation, and $\mathrm{BCR}\in[0,1]$ denotes the boundary coverage ratio of the synthesized vicinal mass.
Interpretation. A higher BCR (i.e., more boundary samples) leads to a lower vicinal risk. If a sufficient fraction of the vicinal mass covers the decision boundary (i.e., large $\mathrm{BCR}$), the loss is Lipschitz-smooth, and off-manifold error $\delta$ is small, then the vicinal risk after augmentation is guaranteed to decrease— by at least $L\,\gamma_{0}\,(\mathrm{BCR}-\zeta)$, up to a small error $\mathcal{O}(\delta)$.
3.3. Graph as a Natural Filter
Here, we outline how the graph structure helps to denoise LLM-based VRM. We introduce a confidence function to assign edges between synthetic and original nodes, ensuring reliable anchoring while isolating noisy samples. Theoretical analysis shows that high-confidence edges guide synthetic nodes toward the correct class manifold, while low-confidence samples remain isolated. Complete proofs of all theorems are provided in Appendix D.
Notation
Let $\bar{\mathcal{V}}$ denote the set of synthetic nodes, and let $\bar{\mathcal{T}}=\{\hat{t}_{\hat{v}}:\hat{v}\in\bar{\mathcal{V}}\}$ be their corresponding texts, where each $\hat{v}$ is a newly generated text-attributed node produced by LLM-based Vicinal Risk Minimization (see $\S$ 3.2). We employ a frozen encoder $\phi:\mathcal{T}\cup\bar{\mathcal{T}}\to\mathcal{X}^{\prime}$ to obtain representations for all texts.
3.3.1. Confidence function.
Using the frozen encoder $\phi$, we map each node $v$ (real or synthetic) to a predictor-space embedding via an MLP encoder $\mathrm{enc}$:
$z_{v}:=\mathrm{enc}\big(\phi(t_{v})\big)\in\mathbb{R}^{p}.$
For any candidate edge $(u,v)$, a second MLP (the predictor, $\mathrm{pred}$) maps their inner-product similarity to a logit:
$s_{u,v}:=\mathrm{pred}\big(\langle z_{u},z_{v}\rangle\big)\in\mathbb{R}.$
We train $\mathrm{enc}$ and $\mathrm{pred}$ on the original graph to score true edges higher than non-edges using a LogSigmoid loss on $s_{u,v}$. At test time, we convert the logit to a calibrated pairwise confidence:
$\kappa(u,v):=\sigma(s_{u,v})\in[0,1],\qquad\sigma(a)=\frac{1}{1+e^{-a}}.$
Ranking by $\kappa$ or by $s$ is equivalent, as $\sigma$ is monotonic.
3.3.2. Confidence-aware edge assignment.